Terephthalic acid is an important chemical used in the manufacture of polyesters and plasticizers. It contains only C, H, and O. Combustion of 19.81 mg terephthalic acid produces 41.98 mg CO2 and 6.45 H2O. If 0.250 mole of terephthalic acid has a mass of 41.5 g, determine the molecular formula for terephthalic acid.
1
Expert's answer
2019-08-29T06:44:00-0400

C8H6O4 it is the answer

Solution

Molecular weigh = 41.5/ 0.25 =166 g/mole

CxHyOz combustion to CO2 and H2O

Mole of CO2 = 41.98*10^-3/44= 0.00095

Mole of H2O = 6.95*10^-3/18= 0.00036

Mole of CxHyOz = 0.00012

Because Cx -> CO2 and

Hy-> H2O, that is why,

X= 0.00095/0.00012= 8

Y= 0.00036*2/0.00012= 6

Z= (166- 8*12- 6*1)/16=4